MICROWAVE-MICROWAVE DOUBLE RESONANCE INVESTIGATION OF THE H_{2}-N_{2}O VAN DER WAALS COMPLEX

\underline{JEN NICOLE LANDRY}, WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, T6G 2G2, Canada ; YOSHIHIRO SUMIYOSHI AND YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguroku, Tokyo 153-8902, Japan.

The J=2-1 rotational transitions of orthoD_{2}-^{15}N_{2}O and paraD_{2}-^{15}N_{2}O were measured using a microwave-microwave double resonance technique. The experimental setup included a pulsed molecular beam Fourier Transform microwave spectrometer for signal detection and a microwave synthesizer with a K band standard gain horn (frequency range 18 - 26.5 GHz) to generate and transmit the pump radiation. The experimental arrangement was such that the pump radiation propagation was perpendicular to the supersonic molecular expansion. In these experiments, the pump radiation was tuned into resonance with the J_{KaKc}=1_{01}-0_{00} or 1_{11}-0_{00} transition. The resulting population transfer from the rotational ground state facilitated the observation of the J=2-1 transition by significantly enhancing the signal intensity. In addition, a brief update on the solvation studies of nitrous oxide with paraH_{2} molecules will be given.